OpenStudy (anonymous):

how do you take the integral of (lnx)/(x^2)?

OpenStudy (anonymous):

INTlnx* x^-2 try integration by parts

OpenStudy (anonymous):

What should my u and v be?

OpenStudy (anonymous):

try u = lnx so du = 1/x and dv = x^-2 so v = - x ^-1 or -1/x

OpenStudy (anonymous):

lets put lnx as u so it'll be int ( u/x du)

OpenStudy (anonymous):

INT = (-1/x)* lnx - INT(-1/x * 1/x dx = (-1/x) ln x + INT x^-2 dx = -1/x) ln x - 1/x = -1/x (ln x+ 1) + c